# Predictability of viral load dynamics in the early phases of SARS-CoV-2 through a model-based approach

### v2

@@ Abstract
Keywords: Mathematical epidemiology; SARS-CoV-2 viral load; SIR with age of infection;
                  Uncertainty quantification methods.
AMS Subject Classification: 92C60; 92C50; 45K05; 65R20; 65C05.
@@ 2.1 Samples and SARS-CoV-2 kinetics
[…]
@@
[…]
@@ 2.3 Data description
[…]
@@ 3 Inference of viral load dynamics
Table 2: Timeline of the main NPIs employed in Italy before the mass vaccination campaign started in mid 2021. The average recovery time from the infection was $14$ days.
@@
over the time horizon $[t_{0},t_{f}]$ representing any of the three periods that can be identified with the initial spreading phase of the three epidemic waves under study: February 24–March 9 2020, October 7–October 22 2020 and February 24–March 10 2021. In particular, we choose here the infection rate $\beta(\tau)$ to be the Gamma function (2) obtained in the previous section from the viral loads measurements, rescaled by the factor $\alpha_{0}=10^{7}$. We use the fitting from November–December 2020 as initial guess for the optimization (6) during the first two epidemic waves, while we opt for the January–May 2021’s fitting to study the third wave. Notice that at this stage the parameters $\alpha_{2}$ and $\alpha_{3}$ are fixed from these two fits, since we only need to ensure a proper initial calibration of the time shift $\alpha_{1}$ to correctly align our model with the epidemiological data. Finally, we always consider a constant recovery rate $\bar{\gamma}=1/14$ to lower the complexity of the minimization. This value is consistent with the existing literature on the COVID-19 infection [14, 33, 51], for which the time to viral clearance has been reported to span approximately between 10 and 21 days on average over the different variants. In Table 3 we collect all the parameters obtained from this initial analysis, for each of the three waves.
@@
At last, in Figures 5 and 6 on the left, we provide a direct comparison between the optimal infection rates from different epidemic waves, obtained via the previous two-step analysis, together with their $80\%$ prediction bounds. These are aligned by their respective peak of infectiousness (which corresponds to day 00 on the graph). One can observe that the infection’s peak progressively decreases, while the tails of the infection rate tend to grow. This suggests that the virus becomes weaker but more effective over a longer period of time and individuals stay infective for more days on average. Moreover, the decrease in curve’s growth in the beginning of the infection could suggest a slower progression of the symptoms and explain the greater spreading of the epidemic with each new wave.

### v3

@@ 2.1 Samples and SARS-CoV-2 kinetics
Individuals that resulted positive from SARS-CoV-2 RNA may leave isolation only when two consecutive negative respiratory samples occurred. Thus, a series of follow-up samples were analyzed to evaluate the structure of SARS-CoV-2 viral loads and used to monitor the evolution of SARS-CoV-2 VL kinetics. For this reason, the VL kinetics of each considered patient are described by at least two samples, one accounting for the time of testing and the other for the measured viral load. In more detail, the experimental data for each patient consist of two $n$-tuples, $n\in\mathbb{N}$, where $n$ is the number of successive molecular tests taken by the individual during the monitoring period: the first $n$-tuple contains the times (in days) when respiratory samples were collected and analyzed (where 0 always represents the day of the first test) and the second contains the kinetics Ct detected by each of these tests (where $40\ \textrm{Ct}$ means complete negativity to the virus and thus recovery). So, for example, a patient that took three successive molecular tests, the first being strongly positive ($\sim 20\ \textrm{Ct}$), the second barely positive ($\sim 36\ \textrm{Ct}$) after 9 days and finally the third negative after 7 more days, would be characterized by the two triples $(0,9,16)$ for the times and $(20,36,40)$ for the kinetics. The values of the corresponding viral loads in terms of RNA $\textrm{copies}/\textrm{ml}$ of respiratory samples may then be deduced from the kinetics of cycle threshold (Ct) using the conversion formula for the quantification of copy number in clinical samples proposed in [40, 41], which is based on the following proportion
@@
where $16.6$ is the dilution factor needed to report VL as RNA $\textrm{copies}/\textrm{ml}$ of respiratory samples. Laboratory data were anonymized and retrospectively analyzed. The study was conducted in accordance with the Declaration of Helsinki and no clinical information were available except for the age and gender.
@@ 2.3 Data description
The study analyzed a total of 233 VL kinetics, corresponding to approximately 700 respiratory samples, collected from patients across two distinct periods. The first period goes from November to December 2020, before circulation of any variants of concern (VOCs), and included 71 patients. The second period, going from January to May 2021 during the circulation of VOCs, involved 162 patients. The subjects considered had a mean age of $42.3\pm 14.7$ years, with $55\%$ of females (120) and $45\%$ of males (113). The demographic distribution, in terms of age and gender, does not statistically differ between the two periods. Using mixed-effects models we found that age and gender did not significantly influence VL kinetics. Time to negativization has a significant effect, showing a decreasing pattern of VL kinetics ($\textrm{estimate}=-482193.7$, $\textrm{p-value}<0.001$). Additionally, a significant difference in VL kinetics was observed between the two periods, with lower values recorded in 2021 compared to 2020 ($\textrm{estimate}=-2122031.9$, $\textrm{p-value}=0.015$).
@@ 3 Inference of viral load dynamics
Table 2: Timeline of the main NPIs employed in Italy before the mass vaccination campaign started in mid 2021. The average recovery time from the infection was 1414 days.
@@
over the time horizon $t_{0},t_{f}]$ representing any of the three periods that can be identified with the initial spreading phase of the three epidemic waves under study: February 24–March 9 2020, October 7–October 22 2020 and February 24–March 10 2021. In particular, we choose here the infection rate $\beta(\tau)$ to be the Gamma function (2) obtained in the previous section from the viral loads measurements, rescaled by the factor $\alpha_{0}=10^{7}$. We use the fitting from November–December 2020 as initial guess for the optimization (6) during the first two epidemic waves, while we opt for the January–May 2021’s fitting to study the third wave. Notice that at this stage the parameters $\alpha_{2}$ and $\alpha_{3}$ are fixed from these two fits, since we only need to ensure a proper initial calibration of the time shift $\alpha_{1}$ to correctly align our model with the epidemiological data. Finally, we always consider a constant recovery rate $\bar{\gamma}=1/14$ to lower the complexity of the minimization. This value is consistent with the existing literature on the COVID-19 infection [14, 33, 51], for which the time to viral clearance has been reported to span approximately between 10 and 21 days on average over the different variants. In Table 3 we collect all the parameters obtained from this initial analysis, for each of the three waves.
@@
At last, in Figures 5 and 6 on the left, we provide a direct comparison between the optimal infection rates from different epidemic waves, obtained via the previous two-step analysis, together with their $80\%$ prediction bounds. These are aligned by their respective peak of infectiousness (which corresponds to day 0 on the graph). One can observe that the infection’s peak progressively decreases, while the tails of the infection rate tend to grow. This suggests that the virus becomes weaker but more effective over a longer period of time and individuals stay infective for more days on average. Moreover, the decrease in curve’s growth in the beginning of the infection could suggest a slower progression of the symptoms and explain the greater spreading of the epidemic with each new wave.
@@ Acknowledgements
Disclaimer
Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or of the European Research Executive Agency (REA). Neither the European Union nor the granting authority can be held responsible for them.